vaporeon246:

Please Help Algebra 2

vaporeon246:

Use \[\log_{12} 3 \approx 0.6826 and \log_{12} 7 \approx 0.7831 \to approximate the value of each exression \]

vaporeon246:

4) \[\log_{12} 36 \] 5) \[\log_{12} 63 \] 6) \[\log_{12} 27/49 \] 9) \[\log_{12} 441 \]

Ultrilliam:

All our math helpers seem to be burried under schoolwork right now except for @dude who is probably busy for the next 30 minutes to a hour and @563blackghost who is AFK ATM x.x

Shadow:

Can you solve the second one?

Shadow:

@vaporeon246

vaporeon246:

i can try

Shadow:

You understand the concept of these problems, right? It's just using the properties of logarithms to break a log down to logs that we know the values of.

vaporeon246:

yes i do @Shadow

Shadow:

\[\log _{12} \frac{ 7 }{ 3 } = \log _{12} 7 - \log _{12} 3\]

vaporeon246:

I was thinking that but i thought it would be wrong

Shadow:

Yeah, these a pretty straightforward if you understand logs. If they are still a touchy topic for you, try view this: https://www.khanacademy.org/math/algebra2/exponential-and-logarithmic-functions/properties-of-logarithms/v/introduction-to-logarithm-properties As Jay said, many of us are busy today. Got two AP quizzes next period c;

Shadow:

Would help if I had the time @vaporeon246

vaporeon246:

That's ok @Shadow

vaporeon246:

here's a somewhat better picture https://mail.google.com/mail/u/0/#inbox/1618c588e3fa17c6